Titanium is manufactured by heating titanium(iv) chloride with sodium. TiCl4(g) + 4Na(l) ->Ti(s) + 4NaCl(s) 

What mass of sodium is required to produce 1 tonne of titanium




1
Expert's answer
2021-05-31T02:25:23-0400

Step I:Calculate the molar mass of Sodium and titanium(iv)

Na=23×2 =46g

Ti=48g


Step II:Write a proportion

46g of Na=48g

x ? =1,000,000g of Ti


"=\frac {1,000,000g×46g}{48g}=958.33Kg"